OpenStudy (anonymous):

Use a transformation to solve the equation. m/-3 = 15 A. m = –45 B. m = –5 C. m = 5 D. m = 45

OpenStudy (agreene):

\[\frac{m}{-3} = 15\] \[-\frac31 \times \frac{m}{-3} = 15 \times -\frac31\] \[\cancel{-\frac31} \times \frac{m}{\cancel{-3}} = 15 \times -3\] \[m = 15 \times -3\]

OpenStudy (anonymous):

thank you